November 1997 Archives

NASA MIR Program / Mir 24 / NASA 6 Status
Date: Thursday, November 20, 1997 // MD 107/56

Mir Station System/Subsystem Activities

Today's major activities centered on preparation and execution of the Cooperative Solar Array experiment. The test consisted of disconnecting, measuring, and connecting multiple solar cells on the Cooperative Solar array (Kvant -Z axis) and the disconnection and re-connection of the Russian Solar array (Kvant +Z axis - installed during the November EVA). Although the Base Block arrays remained connected during the entire test, they were shadowed by Kvant arrays due to the Beta angle and attitude of the experiment. As a precaution, Dave was told to power off IPCG and Glovebox during the comm pass at 18:15 DMT. Additionally, Dave was told to power off PUP A and PUP B for the night passes from 19:00 until 00:15.

At the most critical time of the test, the power margin turned out to be about 20 amps. Mr. Blagov expressed his thanks to the Phase I Program for agreeing to power down some of the science equipment and to cycle the PUP's power.

The only other item of interest we are tracking is the temp in the Base Block. The crew called down temps of 36 degrees (18:16 DMT) and 30 degrees (22:30 DMT). The ground team is attributing the high temperatures to off nominal performance of the Base Block cooling loop KOB-1. The ground is also assessing powering the Base Block air conditioner to make the sleep period more comfortable. With the higher temperatures, the Kvant Electron remained off. The crew reported that they used three oxygen canisters to maintain proper ppO2 levels.

At the 00:11 comm pass on Nov 20/21: The cosmonauts were finishing the the reconnections of the all of the cables and closing out of the panels.

The ground instructed them to turn on both condensate removal pumps. They did and reported that they were operational, even the #2 pump which surprised them. Apparently, there were problems with this pump earlier. Later in the pass, the ground reported that the condensate pumps were not operational per the tm. The ground told the crew this can wait 'til tomorrow.

The ground also instructed the cosmonauts to burn one more candle which gives a total of 3 O2 candles today.

The crew reported that the solar array tracking control was reinitiated. Later in the pass, the crew reported that the solar array drives were nominal

Before the comm pass the Power flight controllers were asking for last minute changes to the attitude, which the SFD took objection to, but gave in. The Mir was in LVLH attitude and the power flight controller wanted to place the station into solar inertial. The fact that the Mir was in LVLH was surprise to us, so Joel has added this to his list of topics for discussion with Yuri Antoshechkin.

Date: Wednesday, November 19, 1997// MD 106/55

Mir Station System/Subsystem Activities

Health of the major systems continue to remain unchanged (Kvant Electron working (periodically turned off/on for power), 11 gyrodynes holding attitude, Condensate System working, Urine Recycling System working).

After continuing discussions with different Russians on last Friday's solar array test, we have learned the experiment was performed to test the Cooperative Solar Array and NOT the newly installed Kvant solar array. Additionally, we learned that the disconnection of the Base Block arrays and the newly installed solar array were not part of this test.

This morning, Victor Blagov told us he attributes the experiment failure to an errant disconnection of one of the Cooperative Solar array connectors. Victor stated this experiment was performed to support the Contract (per Jim Nise, the Contract item is 3FA14). Victor also told us the experiment is scheduled to be repeated Thursday afternoon (after 16:00 DMT) to accomplish the contract requirement.

Today, we also received clarification on the sequence used to bring the gyrodynes back on-line. The Kvant-2 gyrodynes were spun down nominally since power to the module was not interrupted. The spinup of these gyrodynes occurred nominally. However, after the gyrodynes were integrated into the control loop, an out of limit signal was sensed by the precession unit which resulted in two gyrodynes (#2 and #6) spinning down. The out of limit signal was a result of excessive torques when controlling the Mir stack with 6 gyrodynes in one module. These two gyrodynes were spun up after one of the Kvant gyrodynes were back on line and continue to operate nominally.

The Kvant gyrodynes were not spun down nominally since their power is supplied from the Base Block. When testing the gyrodynes for start-up, the ground was unsuccessful in gaining an indication of health status. After two failed attempts to test the health of the gyrodynes, the ground team gave the crew a GO to spin up one gyrodyne (Kvant #4). The gyrodyne spun up nominally and the crew was directed to bring the remaining gyrodynes on-line. After discussing the situation with the motion control experts, the start-up test requires an inhibit command to be lifted prior to the testing. It is believed this inhibit was not released prior to the test.

Prior to the motion control failure, the crew had been working on the Electron unit in Kvant-2 which failed to run due to the O2 sensor in the H2 vent line sensing elevated pressures. The ground controllers believe the sensor may be clogged and have instructed the crew to remove and clean the sensor. These activities will be postponed until late this week or early next week.

Consumables used during the motion control recovery consist of 5 Oxygen canisters, 1 American LiOH canister, approximately 25 kg of Mir propellant, and approximately 10 kg of Soyuz prop.

In addition to the motion control recovery procedures, this week's systems work has focused on preparation and installation of the Base Block Vozdukh. The majority of the crew's time has been to clear out the space where the Vozdukh will be installed. The crew has had multiple questions concerning which hardware and cables were acceptable to move and are concerned that the allotted space will not be sufficient. Vozdukh activities are scheduled for 4-5 hrs/day for the rest of the week.

The ground is watching a decreasing efficiency on the air conditioner compressor. Since a spare is available on-board, the crew will be instructed to use the system as normal until the compressor fails. At that time, the compressor will be replaced with the on-board spare.

Tues evening tm data indicated the ShSO pressure at 278 mmHg.

At the last communication session of the day the crew reported problems with installing the Vozdukh unit in the Base Block. The crew said that there are some compatibility problems with existing pipes on board and that the specialist will have to be consulted tomorrow.


Loading

 



Monthly Archives

About this Archive

This page is an archive of entries from November 1997 listed from newest to oldest.

October 1997 is the previous archive.

December 1997 is the next archive.

Find recent content on the main index or look in the archives to find all content.